OpenStudy (anonymous):

what is the value of p in x= -1/8y^2

OpenStudy (anonymous):

I don't see any p's in that equation... Can you be a bit more specific in what you want answered?

OpenStudy (anonymous):

we are dealing with parabolas and that is all the question says...my teacher says that in a normal parabola it is y= 1/4p x^2

jimthompson5910 (jim_thompson5910):

According to this page http://www.mathwords.com/f/focus_parabola.htm It states that the general form for parabolas that open left/right is 4p(x-h) = (y-k)^2 where p is the focal distance and (h,k) is the vertex. --------------------------------------------------------- So we need to get x= -1/8y^2 into that form x= -1/8y^2 -8x= y^2 ... Multiply both sides by -8. -8(x-0) = (y-0)^2 -4*(2)(x-0) = (y-0)^2 So p = 2 which means the focal distance is 2 units

OpenStudy (anonymous):

so then this would be false the focus of x = -1/8y^2 is (-2,0)

jimthompson5910 (jim_thompson5910):

p = 2 is just the distance the focus is away from the center

jimthompson5910 (jim_thompson5910):

because x= -1/8y^2 opens to the left, we start at the vertex (0,0) then move 2 units along the axis of symmetry y = 0 to land on the focus of (-2,0) So it's actually true

OpenStudy (anonymous):

oh ok that is making more sense now thanks can you help with some more math problems with midterm review?

jimthompson5910 (jim_thompson5910):

Sure I can do a few more

OpenStudy (anonymous):

ok solve the equation for x round to nearest thousandth \[10^{x+5} -2\]

OpenStudy (anonymous):

wait that was a typo hold on

OpenStudy (anonymous):

\[10= 3^{x+5} -2\]

OpenStudy (anonymous):

would i add 2 to 10 and then divide by 3 giving us 4=x+5?

jimthompson5910 (jim_thompson5910):

You do add 2 to both sides, but you cannot divide both sides by 3

OpenStudy (anonymous):

ok then what would i do with the 3?

jimthompson5910 (jim_thompson5910):

We now have \[\Large 3^{x+5} = 12\] we then convert to logarithmic form to get \[\Large x+5 = \log_{3}(12)\] and then subtract 5 from both sides to isolate x \[\Large x = \log_{3}(12)-5\]

jimthompson5910 (jim_thompson5910):

That last equation is the exact answer. Use a calculator to evaluate that to get the approximate answer (make sure you use the correct base)

jimthompson5910 (jim_thompson5910):

You will use the change of base formula \[\Large \log_{b}(x) = \frac{\log(x)}{\log(b)}\]

OpenStudy (anonymous):

i am confused with what to plug into that equation what is log base b of x

jimthompson5910 (jim_thompson5910):

well the last thing I posted is just a general form In this case, b = 3 (base is 3), x = 12 So.... \[\Large \log_{b}(x) = \frac{\log(x)}{\log(b)}\] turns into \[\Large \log_{3}(12) = \frac{\log(12)}{\log(3)}\] after the proper replacements are made

OpenStudy (anonymous):

ok thanks that makes sense

jimthompson5910 (jim_thompson5910):

so what do you get as a final answer?

OpenStudy (anonymous):

2.262

jimthompson5910 (jim_thompson5910):

So that means \[\Large \log_{3}(12) \approx 2.262\] which is correct. Now you must subtract off 5

jimthompson5910 (jim_thompson5910):

Since we found above that \[\Large x = \log_{3}(12)-5\]

OpenStudy (anonymous):

-2.738

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

thanks do u still have time to be on here?

jimthompson5910 (jim_thompson5910):

Yes I do

OpenStudy (anonymous):

mind helping me with more logarthims stuff

jimthompson5910 (jim_thompson5910):

sure I can help, go ahead

OpenStudy (anonymous):

i stink at logarithims lol but what is the range of the funtion show below f(x)= log base 5 (x+2) i believe it would be all real numbers i dont think there are any y restrictions

jimthompson5910 (jim_thompson5910):

You are correct. Nice work.

OpenStudy (anonymous):

thank you now for this one what is domain of f(x)= log base 5 (x-3) i believe it would be greater than or equal to -2

jimthompson5910 (jim_thompson5910):

that is incorrect

jimthompson5910 (jim_thompson5910):

hint: you cannot take the log of a negative number or the log of 0

OpenStudy (anonymous):

all real numbers greater than 3?

jimthompson5910 (jim_thompson5910):

Good, x-3 > 0 x-3+3 > 0+3 x + 0 > 3 x > 3

OpenStudy (anonymous):

thanks so much...i dont wanna keep you any longer. when do you think you will be on here again?

jimthompson5910 (jim_thompson5910):

I should be on later tonight

OpenStudy (anonymous):

it is after 8pm here i am going to try and work on my midterm packet more...maybe you could help me tomorrow or thursday?

jimthompson5910 (jim_thompson5910):

Sure that works for me

OpenStudy (anonymous):

ok sounds good talk to you tomorrow or thursday